Find the equation of a line that is perpendicular to the line y=1/9x+3 ad contains the point (-3,0)
 Answer by lwsshak3(11628)
       
 Find the equation of a line that is perpendicular to the line y=1/9x+3 ad contains the point (-3,0) **
 equation for a straight line:y=mx+b, m=slope, b=y-intercept
 slope of given line=1/9
 slope of line perpendicular to given line=-9 (negative reciprocal)
 equation: y=-9x+b
 solve for b using coordinates of given point (-3,0)
 0=-9*-3+b
 b=-27
 equation: y=-9x-27
 | 
  
 | 
 |